OpenStudy (anonymous):

∫∫1/(x^2+y^2)dxdy solve this converting in polar form limits are x :y to a y :0 to a

OpenStudy (tkhunny):

Trying again? \(\int\limits_0^a\int\limits_y^a \dfrac{1}{x^{2}+y^{2}}\;dxdy = \iint_A \dfrac{r}{r^{2}}\;drd\theta\) You simplfy and tell me the new limits suggested by my "A".

OpenStudy (anonymous):

new limits will be r: 0 to a sec @ @:0 to pie/4

OpenStudy (anonymous):

but now integral will be In r on applying limits I got problem

OpenStudy (tkhunny):

It's a little messy.